Show that < 𝑥𝑝 > − < 𝑝𝑥 > = 𝑖ħ for the ground state wave function of the Quantum 

Harmonic Oscillator (QHO)


1
Expert's answer
2021-02-16T10:28:41-0500

The QHO ground state wave function is given by "\psi(x) = (\frac{m\omega}{\pi \hbar})^{1/4} e^{-\frac{m \omega x^2}{2\hbar}}". In coordinate representation, "\hat{x} = x, \hat{p} =- i\hbar \frac{\partial}{\partial x}" and thus we find :

"<\hat x \hat p >_\psi = \int_{-\infty}^{+\infty} x\cdot (i\hbar\frac{m\omega x}{\hbar})\cdot \psi^2(x) dx", applying the integration by parts ("u = x, dv = im\omega x \psi^2 dx") we find "<\hat x \hat p >_\psi = [\frac{1}{2}x\psi^2]^{+\infty}_{-\infty} + \frac{1}{2}i\hbar \int \psi^2 dx =\frac{ i\hbar}{2}" as "\psi" is a normalised wave function.

Now we calculate "<\hat p \hat x>_\psi = -i\hbar \int (\psi^2+x\psi \psi')dx". The integral "-i\hbar \int \psi^2dx =-i\hbar" is given by the normalisation of "\psi". The second integral was calculated previously. Thus we find "<\hat p \hat x>_\psi = -i\hbar /2" .

Thus "<xp-px>_\psi = i\hbar" which is coherent with the fact that "[x,p]=i\hbar \cdot Id" and thus "<[x,p]>_\psi = i\hbar" for any normalised wave function "\psi".